OpenStudy (unklerhaukus):

The sum of any five consecutive integers is divisible by 5 . Proof:

OpenStudy (unklerhaukus):

Proof: \[\text{If \(x\) is the smallest of the integers,}\]\[\text{the sum of 5 consecutive integers can be written as }\]\[x+(x+1)+(x+2)+(x+3)+(x+4)\]\[\qquad\qquad=5x+(1+2+3+4)\]\[\qquad\qquad=5x+10\]\[\qquad\qquad=5(x+2)\]\[\text{If this is divisible by 5, then }\]\[5|5(x+2)\]\[\text{ is a true statement.}\]\[\text{This is true when \((x+2)\) is an integer which is always }\]\[\text{ true because integers are closed under addition.}\] This proves the sum of any five consecutive integers is divisible by 5. \[\square\]

OpenStudy (unklerhaukus):

is this good?

OpenStudy (anonymous):

Works for me, but I'm pretty lenient when it comes to rigour.

OpenStudy (anonymous):

a + (a+1) + (a+2) + (a+3) + (a+4) = 5a + 10 QED

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):

you dont think i need so many words?

OpenStudy (anonymous):

That's right...and x+2 is neither here nor there....

OpenStudy (anonymous):

5x +10 is already divisible by 5

OpenStudy (anonymous):

Yeah, I do like the extra step of showing that 5x+10 = 5(x+2), then it's obvious that the number is a multiple of five, and hence is divisible by 5.

OpenStudy (anonymous):

The extra step is superfluous (really really superfluous)

jhonyy9 (jhonyy9):

n1=1 n2=2 n3=3 n4=4 n5=5 n1 +n2 +n3 +n4 +n5 =n15 /5 is true n=k suppose is true k1 +k2 +k3 +k4 +k5 =k15 /5 than for k=k+1 (k1 +1) +(k2 +1) +(k3 +1) +(k4 +1) +(k5 +1) = = (k1 +k2 +k3 +k4 +k5) +5= =5+5+5+5= =20/5 Q:E:D:

jhonyy9 (jhonyy9):

oppinions ???

OpenStudy (anonymous):

(a-2) + (a-1) + a + (a+1) + (a+2) = 5a QED

OpenStudy (unklerhaukus):

i like your induction @jhonyy9, but it is long

jhonyy9 (jhonyy9):

so yes thank you but i think that is very very more ,,mathematically" thematically yes???

OpenStudy (unklerhaukus):

my method is a flow of implication of truth

jhonyy9 (jhonyy9):

yes but here on openstudy have told me that a proof need being more mathematically,understandably and not for the last step more understandably for sombody inclusive who not know mathematics just a ,,little" this was for my,, Goldbach's and Collatz"s conjecture proofs"